# On Utility Maximization under Multivariate Fake Stationary Affine Volterra Models

### 11 Mar 2026 v1

@@ 3.1 Optimal strategy for the power utility maximization problem
$=X^{\pi}_{t}\bigl(r(t)+\pi_{t}^{\top}\sigma(V_{t})\lambda_{t}\bigr)\,\mathrm{d}t+X^{\pi}_{t}\pi_{t}^{\top}\sigma(V_{t})\,\mathrm{d}B_{t}=X_{t}^{\pi}(r(t)+\pi_{t}^{\top}{\operatorname{diag}(V_{t})}{\theta})dt+X_{t}\pi_{t}^{\top}\sqrt{\operatorname{diag}(V_{t})}dB_{t},$
@@
$dX^{\alpha}_{t}=\big(r(t)+\alpha_{t}^{\top}\lambda_{t}\big)X^{\alpha}_{t}dt+\alpha_{t}^{\top}X^{\alpha}_{t}dB_{t},\;X_{0}=x_{0}>0.$ (3.30)
@@ 3. $(c)$
where $\mathbb{E}_{x_{0},V_{0}}$ is the conditional expectation given $x_{0}$ and $V_{0}$. The parameter $\gamma$ represents the relative risk aversion of the investor. Smaller $\gamma$ correspond to higher risk aversion. A strategy $\alpha^{*}$ for which the supremum is attained is called an optimal strategy.
@@ 3.1.1 The degenerate correlation case
$J_{t}^{\alpha}=\frac{(X_{t}^{\alpha})^{\gamma}}{\gamma}\Big(\mathbb{E}^{\tilde{\mathbb{P}}}\Big[\operatorname{exp}\Big(\int_{t}^{T}\frac{\gamma}{\delta}\Big(r_{s}+\frac{\left|\lambda_{s}\right|^{2}}{2(1-\gamma)}\Big)ds\Big)|\mathcal{F}_{t}\Big]\Big)^{\delta}=:\frac{(X_{t}^{\alpha})^{\gamma}}{\gamma}\Gamma_{t}\quad\text{where}\quad\delta:=\frac{1-\gamma}{1-\gamma+\gamma\rho^{2}}.$
@@
Note that Equation (5.1) is known in the litterature as a Ricatti backward stochastic differential equation (see e.g. (Abi Jaber et al., 2021, Theorem 3.1) or (Chiu and Wong, 2014, heorem 3.1), upon setting $\tilde{\Lambda}_{t}=\Gamma_{t}\Lambda_{t}$ ).
Existence and uniqueness of the solution to (3.34)-(3.35) are established in (Gnabeyeu, 2026, Theorem A.1) based on the results of Gnabeyeu et al. (b) (see also the Remark on Proposition 3.6).
By considering $\Gamma_{t}$, we overcome the non-Markovian and non-semimartingale difficulty in the variance process (2.2). Now we are ready to give the Ansatz for $J^{\alpha}_{t}$. Consider
$J^{\alpha}_{t}=\frac{X^{\gamma}_{t}}{\gamma}\Gamma_{t}.$ (3.38)
The main result we provide for this case is the following:
@@ 3.1.2 The general correlation case: A verification argument
First of all, we point out that for each of the inhomogeneous Riccati-Volterra equations (3.43)- (3.44) and (3.34)- (3.35), there exists a unique non-continuable continuous solution $(\psi,T_{\operatorname{max}})$ under Assumption 2.1 by (Gnabeyeu, 2026, Theorem A.1) as stated in the following Remark.
@@
(Gnabeyeu, 2026, Theorem A.1) guarantees that there exists a unique non-continuable continuous solution $(\psi,T_{\max})$ to Equation (3.74) with $\psi^{i}\in C([0,T_{max}),\mathbb{R}^{d})$ in the sense that $\psi^{i}$ satisfies (3.74) on $0,T_{max})$ with $T_{max}\in(0,T]$ and $\sup_{t<T_{max}}|\psi^{i}(t)|=+\infty$, if $T_{max}<T$. Combining the component-wise solutions, we finally obtain the existence of a unique non-continuable solution $\psi\in C([0,T_{max}),\mathbb{R}^{d})$ of the inhomogeneous Ricatti–Volterra Equation (3.43)- (3.44) (and in particular of (3.34)– (3.35) ).
@@
We will show directly that the corresponding value attained by this portfolio strategy is $\frac{x_{0}^{\gamma}}{\gamma}\Gamma_{0}$ and every other admissible portfolio strategy will not yield a larger value. Then we have the following verification result:
@@
For the sake of space limitation, but without loss of self-containedness, we present a sketch of the proof here.
@@ 3.2 Optimal strategy for the exponential utility maximization problem
$=X^{\pi}_{t}\bigl(r(t)+\pi_{t}^{\top}\sigma(V_{t})\lambda_{t}\bigr)\,\mathrm{d}t+\pi_{t}^{\top}\sigma(V_{t})\,\mathrm{d}B_{t}=X_{t}^{\pi}(r(t)+\pi_{t}^{\top}{\operatorname{diag}(V_{t})}{\theta})dt+\pi_{t}^{\top}\sqrt{\operatorname{diag}(V_{t})}dB_{t},$
@@
$X_{t}=e^{\int^{t}_{0}r(s)ds}\left(x_{0}+\int_{0}^{t}e^{-\int^{s}_{0}r(u)du}\left(\alpha_{s}^{\top}dB_{s}+\alpha_{s}^{\top}\lambda_{s}\mathrm{d}s\right)\right),$ (3.54)
Note that it is sufficient to assume that $\int_{0}^{t}(\left|\lambda_{s}\right|^{2}+\left|\alpha_{s}\right|^{2})\,\mathrm{d}s<+\infty$ almost surely for all $t\geq 0$ in order to construct the stochastic integrals in Equation (3.54). This boundedness condition holds owing to the inequality $|z|^{2}\leq 2e^{|z|},\,\forall z\in\mathbb{R}$, together with the condition (3.27), and the following admissibility assumption, which is consistent with Hu et al. (2005).
@@ 3.2.1 The degenerate correlation case
$J_{t}^{\alpha}=-\frac{1}{\gamma}\exp\big(-\gamma e^{\int_{t}^{T}r(s)ds}X^{\alpha}_{t}\big)\Big(\mathbb{E}^{\tilde{\mathbb{P}}}\Big[\operatorname{exp}\big(-\frac{1-\rho^{2}}{2}\int_{t}^{T}\left|\lambda_{s}\right|^{2}ds\big)|\mathcal{F}_{t}\Big]\Big)^{\frac{1}{1-\rho^{2}}}=:-\frac{1}{\gamma}\exp\big(-\gamma e^{\int_{t}^{T}r(s)ds}X^{\alpha}_{t}\big)\Gamma_{t}.$
@@
$=\;\Gamma_{t}\Big[\frac{1}{2}\big(\left|\lambda_{t}\right|^{2}+\left|\Sigma\Lambda_{t}\right|^{2}\big)dt+\Lambda_{t}^{\top}d\widetilde{W}_{t}\Big],$ (3.62)
Remark Assume that $K$ satisfies the Assumption 2.1. As $1-\rho^{2}\geq 0$, then (Gnabeyeu, 2026, Theorem A.1) provides the existence of a unique global continuous solution on $[0,T]$ to (3.57)– (3.58).
Consider the below Ansatz for $J^{\alpha}_{t}$
$J^{\alpha}_{t}=-\frac{1}{\gamma}\exp\big[-\gamma e^{\int^{T}_{t}r(s)ds}X_{t}^{\alpha}\big]\Gamma_{t}.$ (3.63)
The main result we provide for this case is the following:
@@ 3.2.2 The general correlation case: A verification argument
First, note that if $K_{i}$ satisfies the Assumption 2.1 for $i=1,\ldots,d$. As $1-\rho_{i}^{2}\geq 0$, then (Gnabeyeu, 2026, Theorem A.1) provides the existence of a unique global continuous solution on $[0,T]$ to (3.66)– (3.67). More details are given in the Remark on Proposition 3.11. In particular, when $\rho_{1}=\ldots=\rho_{d}=\rho$, the function $\psi$ coincides with the unique global solution to the Riccati–Volterra equations (3.57)–(3.58). Consequently, to avoid restrictions on the correlation structure, we still use a verification arguments in the spirit of Bäuerle and Li (2013) to solve the optimization problem, thus extending the results obtained in the preveous section to the more general correlation structure.
@@
where $g$ == $(g^{1},\ldots,g^{d})^{\top}$ is given by (2.19) i.e. the $\mathbb{R}^{d}$-valued process $(g_{t}(s))_{t\leq s}$ is defined in (2.19). Then, $\left(\Gamma,\Lambda\right)$ is a $\mathbb{S}^{\infty}_{\mathbb{F}}([0,T],\mathbb{R})\times L^{2}_{\mathbb{F}}([0,T],\mathbb{R}^{d})$-valued solution to the Riccati BSDE (3.71) below.
@@
Remark: Assume that $K$ satisfies the Assumption 2.1. Then (Gnabeyeu, 2026, Theorem A.1) provides the existence of a unique global continuous solution $\psi\in C([0,T],\mathbb{R}^{d})$ to (3.68)– (3.69) (and in particular to (3.57)– (3.58) ) and $\psi<0$ for $t>0$. More precisely, as the matrix $D$ in the drift of the volatility process is a diagonal matrix, i.e. $D=-\operatorname{diag}{(\lambda_{1},\dots,\lambda_{d})}$, for $i=1,\ldots,d,$ by (Gnabeyeu, 2026, Corollary A.2), since $\frac{\theta_{i}^{2}}{2}<0$, $\psi^{i}\in C([0,T],\mathbb{R}_{-})$ is unique global solution to the following Volterra equation
@@
Moreover, it follows in this case that the condition (3.25) can be made more explicit by bounding $\psi$ with respect to the vector $\theta$. Indeed setting for $i=1,\ldots,d,$ $\bar{\lambda}_{i}:=\inf_{t\in[0,T]}\big(\lambda_{i}+\nu_{i}\rho_{i}\theta_{i}\varsigma^{i}(t)\big)=\lambda_{i}+\nu_{i}\rho_{i}\theta_{i}\|\varsigma^{i}\|_{\infty}{\mathbb{1}_{\rho_{i}\leq 0}},$ which is strictly positive, still by (Gnabeyeu, 2026, Corollary A.2), we have:
@@
For the sake of brevity, yet without compromising self-containment, we present a sketch of the proof here.
@@
4 Numerical experiments: The particular case of the fake stationary rough Heston volatility
@@
$c=\begin{pmatrix}0.01\\ 0.03\end{pmatrix},\quad\mu_{0}=\begin{pmatrix}0.2\\ 0.25\end{pmatrix},\;D=\begin{pmatrix}-0.2&0\\ 0&-0.6\end{pmatrix},\quad\Sigma=\begin{pmatrix}-0.7&0\\ 0&-0.55\end{pmatrix},\;\theta=\begin{pmatrix}0.1\\ 0.1\end{pmatrix},\;\nu=\begin{pmatrix}0.4\\ 0.2\end{pmatrix}$.
@@
4.1 About the numerical scheme for the Volterra and fractional Riccati equations
To simulate the Volterra process (2.2)–(2.18), we first rewrite it using the equivalent Wiener–Hopf transform (see, e.g., (Gnabeyeu and Pagès, 2025; Gnabeyeu et al., a, Propositions 2.8 and 2.2)). We then introduce the $f_{\lambda}$-integrated discrete time Euler-Maruyama scheme defined by the below equation on the time grid $t_{k}=t^{n}_{k}=\frac{kT}{n},k=0,\dots,n$ and in the fractional kernel case, namely recursively for $i=1,2$, $\overline{V}^{i,n}_{0}=V^{i,0}$ and for every $k=1,\ldots,n$,
@@ 4.1 About the numerical scheme for the Volterra and fractional Riccati equations
where the stochastic integrals $\left(I^{i,\ell}_{k}=\int_{t_{\ell-1}}^{t_{\ell}}f_{\alpha_{i},\lambda_{i}}(t_{k}-s)\,dW_{s}\right)$ can be simulated on the discrete grid $(t^{n}_{k})_{0\leq k\leq n}$ by generating an independent sequence of gaussian vectors $G^{n,\ell},\ell=1\cdots n$ using an extended and stable version of Cholesky decomposition of a well-defined covariance matrix $C$. The reader is referred to (Gnabeyeu and Pagès, 2025, Appendix A) for further details about the simulation of the Gaussian stochastic integrals terms in the semi-integrated Euler scheme introduced in this context for Equation (2.2)– (2.18).
To design an approximation scheme for solving the two-dimensional Riccati–Volterra system of equations (3.43)–(3.44) and (3.68)–(3.69) numerically, we employ as in (El Euch and Rosenbaum, 2019, section 5.1) the generalized Adams–Bashforth–Moulton method, often referred to as the fractional Adams method investigated in Diethelm et al. (2002, 2004) as a useful numerical algorithm for solving a fractional ordinary differential equation (ODE) based on a predictor-corrector approach.
Over a regular or uniform discrete time-grid $(t_{k})_{k=1,\ldots,n}$ with mesh or step length $\Delta=\frac{T}{n}$ $(t_{k}=k\Delta)$ for some integer $n$, let $y_{j}^{i}:=\psi^{i}(t_{j})$, the explicit numerical scheme to estimate $\psi$ is given by
$\begin{cases}y_{k+1}^{i,P}=\displaystyle\sum_{j=0}^{k}b^{i}_{j,k+1}f_{i}(t_{j},y_{j}^{i}),\;\text{with}\;f_{i}(t_{j},x):=a_{i}+F_{i}(T-t_{j},x)\\ y_{k+1}^{i}=\displaystyle\sum_{j=0}^{k}a^{i}_{j,k+1}f_{i}(t_{j},y_{j}^{i})+a^{i}_{k+1,k+1}f_{i}(t_{k+1},y_{k+1}^{i,P}),\quad y_{0}^{i}=0\end{cases},\;i\in\{1,2\}.$
where $a_{i}=\frac{\gamma\theta_{i}^{2}}{2(1-\gamma)}$ in section 3.1 (resp. $a_{i}=-\frac{\theta_{i}^{2}}{2}$ in section 3.2), $F_{i}$ in (3.44) (resp. in (3.69)) and the weights $a^{i}_{j,k+1}$, $b^{i}_{j,k+1}$ are defined as
$a^{i}_{j,k+1}=\frac{\Delta^{\alpha_{i}}}{\Gamma(\alpha_{i}+2)}\begin{cases}k^{\alpha_{i}+1}-(k-\alpha_{i})(k+1)^{\alpha_{i}},&\text{if }j=0,\\ (k-j+2)^{\alpha_{i}+1}+(k-j)^{\alpha_{i}+1}-2(k-j+1)^{\alpha_{i}+1},&1\leq j\leq k,\\ 1,&j=k+1,\end{cases}$
and
$b^{i}_{j,k+1}=\frac{\Delta^{\alpha_{i}}}{\Gamma(\alpha_{i}+1)}\left((k+1-j)^{\alpha_{i}}-(k-j)^{\alpha_{i}}\right),\qquad 0\leq j\leq k.$
Here, $T$ denotes the terminal time or time horizon, $n$ the number of time steps, and $\Delta:=\frac{T}{N}$ the time increment.
Theoretical guarantees for the convergence of this numerical algorithm (the 2-dimensional fractional Adams–Bashforth–Moulton method) are established in Li and Tao (2009).
4.2 Numerical illustrations
Figure 1: Graph of the stabilizer $t\to\varsigma_{\alpha_{1},\lambda_{1},c_{1}}(t)$ (left) and 3030 samples paths $t_{k}\mapsto V^{1}_{t_{k}}$ (right) over the time interval $[0,1]$, for the Hurst esponent $H=0.4$, $c_{1}=0.01$ and number of time steps $n=600$.
@@ 4.2 Numerical illustrations
(A) $\gamma=0.2$
(B) $\gamma=0.5$
(C) $\gamma=0.8$
Figure 4: Evolution of the optimal portfolio strategy for different levels of the risk aversion parameter $\gamma$ for both the Power (left) and Exponential (right) utilities functions.
Plots show that, in addition to the impact of the roughness of asset volatility on the optimal allocation documented in Han and Wong (2020); Aichinger and Desmettre (2021), the stabilizing function also affects the optimal portfolio allocation, and hence the hedging demand over time.

### 06 Apr 2026 v2

@@ 3.1 Optimal strategy for the power utility maximization problem
$=X^{\pi}_{t}\bigl(r(t)+\pi_{t}^{\top}\sigma(V_{t})\lambda_{t}\bigr)\,\mathrm{d}t+X^{\pi}_{t}\pi_{t}^{\top}\sigma(V_{t})\,\mathrm{d}B_{t}=X_{t}^{\pi}(r(t)+\pi_{t}^{\top}{\operatorname{diag}(V_{t})}{\theta})dt+X_{t}\pi_{t}^{\top}\sqrt{\operatorname{diag}(V_{t})}dB_{t}.$
@@
$dX^{\alpha}_{t}=\big(r(t)+\alpha_{t}^{\top}\lambda_{t}\big)X^{\alpha}_{t}dt+\alpha_{t}^{\top}X^{\alpha}_{t}dB_{t},\;X_{0}=x_{0}>0.\vskip-2.84544pt$ (3.30)
@@ 3. $(c)$
where $\mathbb{E}_{x_{0},V_{0}}$ is the conditional expectation given $x_{0}$ and $V_{0}$. The parameter $\gamma$ represents the relative risk aversion of the investor. Smaller $\gamma$ correspond to higher risk aversion.
@@ 3.1.1 The degenerate correlation case
$J_{t}^{\alpha}=\frac{(X_{t}^{\alpha})^{\gamma}}{\gamma}\Big(\mathbb{E}^{\tilde{\mathbb{P}}}\Big[\operatorname{exp}\Big(\int_{t}^{T}\frac{\gamma}{\delta}\Big(r_{s}+\frac{\left|\lambda_{s}\right|^{2}}{2(1-\gamma)}\Big)ds\Big)|\mathcal{F}_{t}\Big]\Big)^{\delta}=:\frac{(X_{t}^{\alpha})^{\gamma}}{\gamma}\Gamma_{t}\;\text{where}\;\delta:=\frac{1-\gamma}{1-\gamma+\gamma\rho^{2}}.$ (3.34)
@@
Note that Equation (A.1) is known in the litterature as a Ricatti backward stochastic differential equation (see e.g. (Abi Jaber et al., 2021, Theorem 3.1) or (Chiu and Wong, 2014, Theorem 3.1), upon setting $\tilde{\Lambda}_{t}=\Gamma_{t}\Lambda_{t}$ ).
Existence and uniqueness of the solution to (3.35)-(3.36) are established in (Gnabeyeu, 2026, Theorem A.2.) based on the results of Gnabeyeu et al. (2026) (see also the Remark on Proposition 3.6).
By considering $\Gamma_{t}$, we overcome the non-Markovian and non-semimartingale difficulty in the variance process (2.2).
Remark: One should note that the Ansatz or the constructed family of stochastic processes $(J^{\alpha})_{\alpha}$ in (3.34) can be rewritten for every $t\in[0,T]$ as
$J^{\alpha}_{t}=\frac{x^{\gamma}_{0}}{\gamma}\exp\!\left(\int_{0}^{t}\gamma\Big(r(s)+\alpha_{s}^{\top}\lambda_{s}-\tfrac{1}{2}\left|\alpha_{s}\right|^{2}\Big)\,ds+\int_{0}^{t}\gamma\alpha_{s}^{\top}\,dB_{s}+Y_{t}\right),\quad x_{0}\geq 0.$ (3.39)
where the pair $(Y,\Lambda)$ satisfies a rather simple backward stochastic differential equation (BSDE) under $\tilde{\mathbb{P}}$ with the final condition $Y_{T}=0$, so that $\Gamma_{t}:=\exp\big(Y_{t}\big)$ satisfies the BSDE (A.1). This approach is analogous to that in (Hu et al., 2005, Equation 13). The main result we present in this case is as follows:
@@ 3.1.2 The general correlation case: A verification argument
First of all, we point out that for each of the inhomogeneous Riccati-Volterra equations (3.44)- (3.45) and (3.35)- (3.36), there exists a unique non-continuable continuous solution $(\psi,T_{\operatorname{max}})$ under Assumption 2.1 by (Gnabeyeu, 2026, Theorem A.2.) as stated in the following Remark.
@@
(Gnabeyeu, 2026, Theorem A.2.) guarantees that there exists a unique non-continuable continuous solution $(\psi,T_{\max}^{i})$ to Equation (3.49) with $\psi^{i}\in C([0,T_{max}^{i}),\mathbb{R})$ in the sense that $\psi^{i}$ satisfies (3.49) on $0,T_{max}^{i})$ with $T_{max}^{i}\in(0,T]$ and $\sup_{t<T_{max}^{i}}|\psi^{i}(t)|=+\infty$, if $T_{max}^{i}<T$. Combining the component-wise solutions, we finally obtain the existence of a unique non-continuable solution $\psi\in C([0,T_{max}),\mathbb{R}^{d})$ of the inhomogeneous Ricatti–Volterra Equation (3.44)- (3.45) (and in particular of (3.35)– (3.36) ).
@@
We will directly show that the value attained by this portfolio strategy is $\frac{x_{0}^{\gamma}}{\gamma}\Gamma_{0}$, and that no other admissible portfolio strategy can yield a larger value, through the following verification result:
@@
For the sake of space limitation, but without loss of self-containedness, we present a sketch of the proof.
@@ 3.2 Optimal strategy for the exponential utility maximization problem
$=X^{\pi}_{t}\bigl(r(t)+\pi_{t}^{\top}\sigma(V_{t})\lambda_{t}\bigr)\,\mathrm{d}t+\pi_{t}^{\top}\sigma(V_{t})\,\mathrm{d}B_{t}=X_{t}^{\pi}(r(t)+\pi_{t}^{\top}{\operatorname{diag}(V_{t})}{\theta})dt+\pi_{t}^{\top}\sqrt{\operatorname{diag}(V_{t})}dB_{t}.$
@@
$X_{t}=e^{\int^{t}_{0}r(s)ds}\Big(x_{0}+\int_{0}^{t}e^{-\int^{s}_{0}r(u)du}\Big(\alpha_{s}^{\top}dB_{s}+\alpha_{s}^{\top}\lambda_{s}\mathrm{d}s\Big)\Big).$ (3.55)
Note that it is sufficient to assume that $\int_{0}^{t}(\left|\lambda_{s}\right|^{2}+\left|\alpha_{s}\right|^{2})\,\mathrm{d}s<+\infty$ almost surely for all $t\geq 0$ in order to construct the stochastic integrals in Equation (3.55). This boundedness condition holds owing to the inequality $|z|^{2}\leq 2e^{|z|},\,\forall z\in\mathbb{R}$, together with the condition (3.27), and the following admissibility assumption, which is consistent with (Hu et al., 2005, Definition 1).
@@ 3.2.1 The degenerate correlation case
$J_{t}^{\alpha}=-\frac{1}{\gamma}\exp\big(-\gamma e^{\int_{t}^{T}r(s)ds}X^{\alpha}_{t}\big)\Big(\mathbb{E}^{\tilde{\mathbb{P}}}\Big[\operatorname{exp}\big(-\frac{1-\rho^{2}}{2}\int_{t}^{T}\left|\lambda_{s}\right|^{2}ds\big)|\mathcal{F}_{t}\Big]\Big)^{\frac{1}{1-\rho^{2}}}=:-\frac{1}{\gamma}\exp\big(-\gamma e^{\int_{t}^{T}r(s)ds}X^{\alpha}_{t}\big)\Gamma_{t}.$ (3.58)
@@
$=\;\Gamma_{t}\Big[\frac{1}{2}\big(\left|\lambda_{t}\right|^{2}+\left|\Sigma\Lambda_{t}\right|^{2}\big)dt+\Lambda_{t}^{\top}d\widetilde{W}_{t}\Big],\quad\Gamma_{T}=1.$ (3.64)
Remark Assume that $K$ satisfies the Assumption 2.1. As $1-\rho^{2}\geq 0$, then (Gnabeyeu, 2026, Theorem A.2.) provides the existence of a unique global continuous solution on $[0,T]$ to (3.59)– (3.60). The main result we provide for this case is the following:
@@ 3.2.2 The general correlation case: A verification argument
First, note that if $K_{i}$ satisfies the Assumption 2.1 for $i=1,\ldots,d$. As $1-\rho_{i}^{2}\geq 0$, then (Gnabeyeu, 2026, Theorem A.2.) provides the existence of a unique global continuous solution on $[0,T]$ to (3.67)– (3.68). More details are given in the Remark on Proposition 3.11. In particular, when $\rho_{1}=\ldots=\rho_{d}=\rho$, the function $\psi$ coincides with the unique global solution to the Riccati–Volterra equations (3.59)–(3.60). Consequently, to avoid restrictions on the correlation structure, we still use a verification arguments in the spirit of Bäuerle and Li (2013) to solve the optimization problem, thus extending the results obtained in the preveous section to the more general correlation structure.
@@
where $g$ == $(g^{1},\ldots,g^{d})^{\top}$ is given by (2.19) i.e. the $\mathbb{R}^{d}$-valued process $(g_{t}(s))_{t\leq s}$ is defined in (2.19).
Then, $\left(\Gamma,\Lambda\right)$ is a $\mathbb{S}^{\infty}_{\mathbb{F}}([0,T],\mathbb{R})\times L^{2}_{\mathbb{F}}([0,T],\mathbb{R}^{d})$-valued solution to the Riccati BSDE (3.72) below.
@@
Remark: Assume that $K$ satisfies the Assumption 2.1. Then (Gnabeyeu, 2026, Theorem A.2.) provides the existence of a unique global continuous solution $\psi\in C([0,T],\mathbb{R}^{d})$ to (3.69)– (3.70) (and in particular to (3.59)– (3.60) ) and $\psi<0$ for $t>0$. More precisely, as the matrix $D$ in the drift of the volatility process is a diagonal matrix, i.e. $D=-\operatorname{diag}{(\lambda_{1},\dots,\lambda_{d})}$, for $i=1,\ldots,d,$ by (Gnabeyeu, 2026, Theorem A.2. $(c)$), since $\frac{\theta_{i}^{2}}{2}<0$, $\psi^{i}\in C([0,T],\mathbb{R}_{-})$ is unique global solution to the following Volterra equation
@@
Moreover, it follows in this case that the condition (3.25) can be made more explicit by bounding $\psi$ with respect to the vector $\theta$. Indeed setting for $i=1,\ldots,d,$ $\bar{\lambda}_{i}:=\inf_{t\in[0,T]}\big(\lambda_{i}+\nu_{i}\rho_{i}\theta_{i}\varsigma^{i}(t)\big)=\lambda_{i}+\nu_{i}\rho_{i}\theta_{i}\|\varsigma^{i}\|_{\infty}{\mathbb{1}_{\rho_{i}\leq 0}},$ (owing to Example 2.6 for the function $\varsigma$) and assuming that $\bar{\lambda}_{i}\neq 0$, by (Gnabeyeu, 2026, Corollary A.3.), we have:
@@
For the sake of brevity, yet without compromising self-containment, we present a sketch of the proof.
@@
3.3 Optimal strategy for the logarithmic utility maximization problem
To complete the spectrum of important utility functions, in this section we shall consider logarithmic utility. Here, we assume the model (2.14)– (2.16) for $S_{t}$, that the dynamics of the controlled wealth process is given by equations (3.30)–(3.31) and the utility function is of the form $U(x):=\log(x)$.
Let the set of all admissible investment strategies, still denoted as $\mathcal{A}$} be given by
[…]
We want to solve the Merton logarithmic utility optimization case, i.e. our aim is to find the value function defined in (3.21) for the logarithmic utility function such that
$\mathcal{V}(x_{0},v_{\infty}):=\operatorname*{\,sup}_{\alpha(\cdot)\in\mathcal{A}}\mathbb{E}\Big[\log(X_{T}^{\alpha})\Big],\qquad X_{0}^{\alpha}:=x_{0}>0,\quad v_{\infty}:=\mathbb{E}[V_{0}];$ (3.83)
Theorem 3.13.
Let $T>0$ be fixed and that Assumption 3.1 is in force. Then for $t\in[0,T]$, an optimal investment strategy $(\alpha_{t}^{*})_{t\in[0,T]}$ for the Merton portfolio problem (3.83) is given by
$\alpha^{*}_{t}$ $=\lambda_{t}=\;\Big(\theta_{i}\sqrt{V_{t}^{i}}\Big)_{1\leq i\leq d},\quad 0\leq t\leq T.$ (3.84)
Moreover $\alpha^{*}$ is admissible and the value function or maximum expected utility defined in (3.83) satisfies
$\mathcal{V}(x_{0},v_{\infty}):=\log(x_{0})+\int_{0}^{T}r(s)ds+\frac{T}{2}\sum_{i=1}^{d}\theta_{i}^{2}\,v_{\infty}^{i},\quad v_{\infty}^{i}:=\mathbb{E}[V_{0}^{i}]=\frac{1-a_{i}}{1-a_{i}\varphi_{\infty}^{i}}\frac{\mu_{\infty}^{i}}{\lambda_{i}}\text{\penalty 10000\ for\penalty 10000\ }i=1,\ldots,d.$ (3.85)
4 Numerical experiments: The fake stationary rough Heston volatility
@@
$c=\begin{pmatrix}0.01\\ 0.03\end{pmatrix},\quad\mu_{0}=\begin{pmatrix}2.0\\ 2.5\end{pmatrix},\;D=\begin{pmatrix}-0.2&0\\ 0&-0.6\end{pmatrix},\quad\Sigma=\begin{pmatrix}-0.7&0\\ 0&-0.55\end{pmatrix},\;\theta=\begin{pmatrix}0.1\\ 0.1\end{pmatrix},\;\nu=\begin{pmatrix}0.4\\ 0.2\end{pmatrix}$.
@@
To simulate the Volterra process (2.2)–(2.18), we first rewrite it using the equivalent Wiener–Hopf transform (see, e.g., (Gnabeyeu and Pagès, 2025; Gnabeyeu et al., 2025, Propositions 2.8 and 2.2)). We then introduce the $f_{\lambda}$-integrated discrete time Euler-Maruyama scheme defined by the below equation on the time grid $t_{k}=t^{n}_{k}=\frac{kT}{n},k=0,\dots,n$ and in the fractional kernel case, namely recursively for $i=1,2$, $\overline{V}^{i,n}_{0}=V^{i,0}$ and for every $k=1,\ldots,n$,
@@
where the stochastic integrals $\left(I^{i,\ell}_{k}=\int_{t_{\ell-1}}^{t_{\ell}}f_{\alpha_{i},\lambda_{i}}(t_{k}-s)\,dW_{s}\right)$ can be simulated on the discrete grid $(t^{n}_{k})_{0\leq k\leq n}$ by generating an independent sequence of gaussian vectors $G^{n,\ell},\ell=1\cdots n$ using an extended and stable version of Cholesky decomposition of a well-defined covariance matrix $C$. The reader is referred to (Gnabeyeu and Pagès, 2025, Appendix A) for further details about the simulation of the Gaussian stochastic integrals terms in the semi-integrated Euler scheme introduced in this context for Equation (2.2)– (2.18). Theoretical guarantees for the convergence of this numerical scheme, as well as the convergence rate are established in Gnabeyeu and Pagès (2026) for more general Kernels and path-dependent coefficients.
To numerically solve the two-dimensional Riccati–Volterra system (3.44)–(3.45) and (3.69)–(3.70), we employ, as in Gnabeyeu (2026) (see also (El Euch and Rosenbaum, 2019, Section 5.1)), the generalized Adams–Bashforth–Moulton scheme, also known as the fractional Adams method. This predictor–corrector method, introduced in Diethelm et al. (2002, 2004) for fractional ordinary differential equations, comes with proven convergence guarantees established in Li and Tao (2009).
Figure 1: Graph of the stabilizer $t\to\varsigma_{\alpha_{1},\lambda_{1},c_{1}}(t)$ (left) and $30$ samples paths $t_{k}\mapsto V^{1}_{t_{k}}$ (right) over the time interval $[0,1]$, for the Hurst esponent $H=0.4$, $c_{1}=0.01$ and number of time steps $n=600$.
@@
The drift and stabilizing functions $\varsigma$ introduced in Proposition 2.4 and Example 2.6 are designed to ensure that the volatility processes $(V^{1},V^{2})$ possess constant marginal means (Figure 1) and variances (Figure 2) over time. This guarantee the invariance of the first two moments under time shifts.
Figures 3 below also confirm the Remark on Proposition 3.6, that is the claim that $\psi\leq 0$ for every $\rho_{i}\in(0,1)$ and every $i\in\{1,2\}$ in the exponential utility case.
@@
(a) $\gamma=0.2$ and $T=1$
(b) $\gamma=0.8$ and $T=5$
Figure 4: Evolution of the optimal portfolio strategy for different levels of the risk aversion parameter $\gamma$ with $T\in\{1,5\}$ for both the Power (left) and Exponential (right) utilities functions.
Plots show that, in addition to the impact of the roughness of asset volatility on the optimal allocation documented in Han and Wong (2020); Aichinger and Desmettre (2021), the stabilizing function also affects the optimal portfolio allocation, and hence the hedging demand over time. The form of the control is consistent across different time scales. The roughness dominates the very short end, while the stabilizing function dictates the long-term behaviour. While the rough volatility models capture high-frequency dynamics with remarkable accuracy, their lack of stationarity make them unsuitable for long-term investment analysis. This framework is then also essential for formulating a well-posed infinite-horizon investment problem, time-series analysis and long-term econometric estimation.
Our illustrations also show that the optimal allocation in the exponential utility case is moderated as the paths of the volatility become rougher. Since $\varsigma^{i}(0)\psi^{i}(T)=\varsigma^{i}(T)\psi^{i}(0)=0$ for $i\in\{1,2\}$, it follows that, in the power utility case, the vector of amounts invested in each stock coincides at the beginning and at the end of the time horizon, as illustrated in the left panels of Figures 4.